A restaurant hands out a scratch-off game ticket with prizes being worth purchases at the restaurant. The back of the ticket lists the odds of winning each dollar value: 0.4 for $5, 0.1 for $10, 0.03 for $75, and 0.005 for $100. What are the odds that the ticket is worth at least $75?
Answer #1

probability that ticket is at least $75 = P(X=75)+P(X=100)

=0.03/(1+0.03)+0.005/(1+0.005)

=0.03410

therefore odds that the ticket is worth at least $75 =probability/(1-probability)=0.03410/(1-0.03410)

=0.03530